I am doing homework the problem is....
A Jar contains 21 nickels and dimes The total value of money inside the jar is $1.70. How many nickels and dimes are there?

I am not understanding how to convert it into 2 different equations.
Thank you

Answer by ankor@dixie-net.com(22740)

let n = no. of nickels
let d = no. dimes
:
Write an equation for each statement
:
A Jar contains 21 nickels and dimes
n + d = 21
n = (21-d); use this form for substitution
:
The total value of money inside the jar is $1.70.
.05n + .10d = 1.70
Substitute (21-d) for n
.05(21-d) + .10d = 1.70
1.05 - .05d + .10d = 1.70\
-.05d + .10d = 1.70 - 1.05
.05d = .65
d = .65/.05
d = 13 dime
I'll let you find the number of nickels
Check your solutions in the 2nd equation






How many nickels and dimes are there?